# BoundMatch: Boundary detection applied to semi-supervised segmentation for urban-driving scenes

### v3

@@ Abstract
Semi-supervised semantic segmentation (SS-SS) aims to mitigate the heavy annotation burden of dense pixel labeling by leveraging abundant unlabeled images alongside a small labeled set. While current consistency regularization methods achieve strong results, they often overlook a critical challenge: the precise delineation of object boundaries. In this paper, we propose BoundMatch, a novel multi-task SS-SS framework that explicitly integrates semantic boundary detection into a teacher-student consistency regularization pipeline. Our core mechanism, Boundary Consistency Regularized Multi-Task Learning (BCRM), enforces prediction agreement between teacher and student models on both segmentation masks and detailed semantic boundaries. To further enhance performance and sharpen boundaries, BoundMatch incorporates two lightweight fusion modules: Boundary-Semantic Fusion (BSF) injects learned boundary cues into the segmentation decoder, while Spatial Gradient Fusion (SGF) refines boundary predictions using mask gradients, leading to higher-quality boundary pseudo-labels. This framework is built upon SAMTH, a strong teacher-student baseline featuring a Harmonious Batch Normalization (HBN) update strategy for improved stability. Extensive experiments on diverse urban-driving scene datasets including Cityscapes, BDD100K, and SYNTHIA show that BoundMatch achieves competitive performance against current state-of-the-art methods. Our approach achieves state-of-the-art results on the new benchmark with DINOv2 foundation model. We further validate our approach’s generalizability on Pascal VOC and ADE20K datasets. Ablation studies highlight BoundMatch’s ability to improve boundary-specific evaluation metrics, its effectiveness in realistic large-scale unlabeled data scenarios, and applicability to lightweight architectures for mobile deployment.
@@ 1 Introduction
Semantic segmentation, the task of assigning a class label to every pixel, provides the detailed scene understanding crucial for many modern vision systems. It is particularly vital for autonomous driving, where precise identification of road boundaries, pedestrians, and vehicles is paramount for safe navigation [1, 2]. Beyond driving, it empowers robotic agents to interact intelligently with complex environments [3, 4], aids in clinical diagnostics through medical image analysis [5, 6], and enables large-scale environmental monitoring via remote sensing [7]. However, the reliance of high-performance models on vast amounts of dense, pixel-level annotations presents a major bottleneck, as manual labeling is extremely time-consuming and costly. To address this issue, semi-supervised semantic segmentation (SS-SS) offers a compelling solution by aiming to achieve strong performance while drastically reducing annotation requirements, primarily leveraging large quantities of readily available unlabeled data.
Among various SS-SS strategies, Consistency Regularization (CR) has proven highly effective and is widely adopted, often implemented within teacher-student frameworks [8, 9]. In this setup, a “teacher” model (e.g. an exponential moving average (EMA) of the “student”) generates pseudo-labels on weakly perturbed unlabeled images, which then supervise the student model trained on strongly perturbed versions of the same images. This encourages the model to produce consistent predictions despite input perturbations. While CR methods have significantly boosted overall segmentation accuracy (e.g. mIoU), their improvements are primarily observed in interior regions, with boundary areas remaining challenging. This limitation hinders the reliability of downstream tasks by adding uncertainty (e.g. a pedestrian merging into the background). Notably, state-of-the-art fully supervised methods explicitly address this issue by incorporating boundary detection as an auxiliary multi-task learning (MTL) objective, yielding demonstrably sharper results [10, 11, 12].
Despite the proven benefits of boundary information in fully supervised settings, its integration into semi-supervised learning remains largely underexplored. Existing attempts fall into two categories: (1) methods that derive boundary cues from segmentation pseudo-labels to re-weight consistency losses near edges [13, 14], and (2) approaches like BoundaryMatch [15] that add binary boundary detection as an auxiliary task. The former relies on potentially noisy pseudo-label boundaries, while the latter treats boundary detection as a disconnected auxiliary task without explicit mechanisms to integrate boundary information back into segmentation, resulting in minimal performance gains. This reveals a crucial gap: the opportunity to enforce CR on semantic boundaries (rather than binary edges) derived from hierarchical features, while creating explicit fusion pathways that allow boundary and segmentation tasks to mutually enhance each other throughout the learning process.
@@
To fill this gap, we propose BoundMatch, a novel teacher-student framework that effectively integrates semantic boundary detection with CR for SS-SS. At its core, BoundMatch introduces Boundary Consistency Regularized Multi-Task Learning (BCRM), which enforces prediction agreement between teacher and student on both segmentation masks and semantic boundaries derived from hierarchical features. To create a synergistic relationship between tasks, we incorporate two fusion modules: Boundary-Semantic Fusion (BSF) injects boundary cues into the segmentation decoder, while Spatial Gradient Fusion (SGF) refines boundary predictions using spatial gradients, yielding higher-quality pseudo-labels. Additionally, we develop SAMTH, a strong teacher-student baseline that addresses training instabilities through Harmonious Batch Normalization (HBN), which maintains consistent batch normalization (BN) statistics between teacher and student networks. BoundMatch is designed as a flexible framework that can enhance various CR methods, as we demonstrate by integrating it with multiple baselines including SAMTH, UniMatch, and PrevMatch.
Extensive experiments demonstrate that BoundMatch achieves competitive or state-of-the-art performance across diverse benchmarks for urban-driving scenes, with particularly notable improvements in boundary quality metrics and strong results when using DINOv2 foundation models. BoundMatch could also generalize to other datasets such as ADE20K and Pascal VOC, and it maintains effectiveness in realistic large-scale unlabeled data scenarios.
@@
  • We provide a new perspective for SS-SS: that semantic boundaries can serve as an effective auxiliary task for consistency regularization. Through BCRM, we demonstrate that enforcing consistency on semantic boundaries (not just binary edges) derived from hierarchical features opens new possibilities for improving semi-supervised segmentation.
  • We propose BoundMatch, a comprehensive framework that realizes this perspective through careful architectural design. Beyond BCRM, BoundMatch incorporates two crucial fusion modules, BSF and SGF, that create bidirectional information flow between boundary and segmentation tasks. This synergistic design enables significant performance gains by allowing both tasks to mutually improve each other’s pseudo-labels.
  • We identify and resolve training instabilities in teacher-student frameworks using EMA update. Our HBN strategy yields SAMTH, a simple yet strong baseline that achieves competitive results even without boundary modeling, demonstrating the importance of proper BN handling in SS-SS.
  • We emphasize the importance of boundary-specific evaluation in SS-SS by systematically reporting BIoU and BF1 metrics alongside traditional mIoU, revealing that boundary quality improvements not captured by standard metrics.
  • Through extensive experiments across five datasets, we validate BoundMatch’s effectiveness when applied to various baselines, achieving state-of-the-art results with DINOv2 backbones, and demonstrating successful integration with lightweight architectures for real-world applications.
@@ 2 Related Work
Fully supervised semantic segmentation methods require large amounts of costly pixel-wise annotations. SS-SS aims to mitigate this burden by leveraging abundant unlabeled data alongside a smaller labeled set. Consistency regularization (CR), which encourages models to produce consistent predictions under different input perturbations, is a cornerstone technique in SS-SS [30, 31, 32, 33, 34, 35, 36]. A common implementation involves the mean-teacher framework [8], where an exponential moving average (EMA) of the student model’s weights generates stable pseudo-labels for unlabeled data. CutMix-MT [8] integrated CutMix augmentation to enhance consistency. Subsequent works like AugSeg [37] and iMAS [38] explored stronger augmentation strategies. Teacher-Student framework has been further extended with various techniques, such as using multiple teachers (e.g. DualTeacher [39]) and contrastive learning (e.g. ReCo [40], U2PL [41]).
There are other notable directions in CR. Some methods focus on co-training to provide cross supervision by using two or more independent networks with different initializations or architectures (e.g. CPS [42] and Diverse Co-Training [43]). On the other hand, approaches like UniMatch [9] and its variants [44, 45, 46] have proposed a unified framework that combines strong-to-weak perturbation regularization with a single model training strategy.
@@
While most SS-SS methods concentrate on pseudo-label refinement and consistency losses, explicit integration of boundary information remains rare. Some approaches derive boundary masks from segmentation pseudo-labels using simple operators: CFCG [13] applies Laplacian operators while CW-BASS [14] uses Sobel filters. These masks are then used to re-weight the cross-entropy loss and put importance on boundary regions. While these methods demonstrate that boundary awareness can improve SS-SS, they are fundamentally limited to the boundary information already present in the pseudo-labels.
More relevant to our work are methods that explicitly incorporate boundary detection as an auxiliary task, learning boundaries independently rather than deriving them from segmentation outputs. Graph-BAS3Net [5] pioneered this direction for semi-supervised medical image segmentation. Most notably, BoundaryMatch [15] recently integrated binary boundary detection into SS-SS by adding an auxiliary boundary head atop UniMatch, following a design similar to STDC [29]. However, BoundaryMatch reported only marginal improvements when integrated with modern SS-SS methods, raising important questions about how to effectively leverage boundary detection in semi-supervised settings.
Our BoundMatch addresses these challenges through three key design decisions that differentiate it from BoundaryMatch: (1) We employ semantic (multi-label) boundaries rather than binary edges, providing richer class-specific supervision that better captures the nuances of different object categories; (2) We leverage hierarchical features from multiple backbone stages [11] instead of single-stage features, enabling multi-scale boundary reasoning; (3) Most critically, we introduce explicit fusion mechanisms (BSF and SGF) that create bidirectional information flow between boundary and segmentation tasks, rather than treating them as independent objectives. These design choices enable BoundMatch to achieve significant performance gains that we comprehensively measure using both standard mIoU and boundary-specific metrics (BIoU and BF1). Our systematic evaluation reveals that boundary-aware SS-SS methods provide substantial improvements in boundary quality that traditional metrics fail to capture, demonstrating that the key to effective boundary-aware SS-SS lies in creating true synergy between the auxiliary and main tasks rather than simply adding boundary detection as a disconnected component.
@@ 3 Approach
In this section, we present BoundMatch, our multi-task framework that incorporates boundary detection into semi-supervised semantic segmentation (SS-SS). We begin by outlining the notation and preliminaries for SS-SS, including the teacher-student paradigm, in Secs. 3.1 and 3.2. We then introduce SAMTH, our baseline method detailed in Sec. 3.3. Building upon this baseline, Sec. 3.4 describes our core Boundary Consistency Regularized Multi-Task Learning (BCRM) framework. Subsequently, Sec. 3.5 details the complementary boundary-aware modeling and refinement strategies: Boundary-Semantic Fusion (BSF) and Spatial Gradient Fusion (SGF). Finally, Sec. 3.6 synthesizes these components (SAMTH, BCRM, BSF, and SGF) into the complete BoundMatch framework.
@@ 3.2 Teacher-Student Framework for Consistency Regularization
The teacher-student framework, as shown in Fig. 2, is a popular paradigm for SS-SS. It employs consistency regularization (CR) between the predictions of a teacher network and a student network for unlabeled samples. The core idea is to encourage the student network to produce consistent predictions for an unlabeled image under different perturbations. The teacher network, which is typically an exponential moving average (EMA) of the student network, provides more stable and reliable predictions, which are then used to guide the student’s learning.
@@ 3.3 Strong Data Augmentation Mean-Teacher with Hard Pseudo-Labels (SAMTH)
HBN aims to mitigate these concerns. We forward the complete batch ($x,u^{w},u^{s}$) through both networks: $p^{S,L},p^{S,w},p^{S,s}=f_{\theta^{S}}(x,u^{w},u^{s})$ and $p^{T,L},p^{T,w},p^{T,s}=f_{\theta^{T}}(x,u^{w},u^{s})$. Crucially, the teacher’s BN layers remain in train mode, updating statistics independently based on the full data stream passing through the teacher’s own parameters. For every iteration, the teacher updates its BN statistics, while the rest of the model parameters (i.e. layer weights) are EMA-updated from the student’s parameters. This ensures the teacher’s BN statistics are harmonious with its specific weights and the full input distribution, decoupling its normalization from student fluctuations. This promotes more stable training dynamics and improves performance, as validated empirically across multiple baselines in our ablations (Sec. 4.3.6). For more complete details and pseudocode, please refer to Appendix A.

### v4

@@ Abstract
Semi-supervised semantic segmentation (SS-SS) aims to mitigate the heavy annotation burden of dense pixel labeling by leveraging abundant unlabeled images alongside a small labeled set. While current consistency regularization methods achieve strong results, most do not explicitly model boundaries as a separate learning objective. In this paper, we propose BoundMatch, a novel multi-task SS‑SS framework that explicitly integrates semantic boundary detection into a teacher-student consistency regularization pipeline. Our core mechanism, Boundary Consistency Regularized Multi-Task Learning (BCRM), enforces prediction agreement between teacher and student models on both segmentation masks and detailed semantic boundaries, providing complementary supervision from two independent tasks. To further enhance performance and encourage sharper boundaries, BoundMatch incorporates two lightweight fusion modules: Boundary-Semantic Fusion (BSF) injects learned boundary cues into the segmentation decoder, while Spatial Gradient Fusion (SGF) refines boundary predictions using mask gradients, yielding more reliable boundary pseudo-labels. This framework is built upon SAMTH, a strong teacher-student baseline featuring a Harmonious Batch Normalization (HBN) update strategy for improved stability. Extensive experiments on diverse datasets including Cityscapes and Pascal VOC show that BoundMatch achieves competitive performance against current state-of-the-art methods. Our approach achieves state-of-the-art results on the new Cityscapes benchmark with DINOv2 foundation model. Ablation studies highlight BoundMatch’s ability to improve boundary-specific evaluation metrics, its effectiveness in realistic large-scale unlabeled data scenario, and applicability to lightweight architectures for mobile deployment.
@@ 1 Introduction
Semantic segmentation, the task of assigning a class label to every pixel, provides the detailed scene understanding crucial for many modern vision systems. It is particularly vital for autonomous driving, where precise identification of road boundaries, pedestrians, and vehicles is paramount for safe navigation [1, 2]. Beyond driving, it empowers robotic agents to interact intelligently with complex environments [3, 4], aids in clinical diagnostics through medical image analysis [5, 6], and enables large-scale environmental monitoring via remote sensing [7]. However, the reliance of high-performance models on vast amounts of dense, pixel-level annotations presents a major bottleneck, as manual labeling is extremely time-consuming and costly. To address this issue, semi‑supervised semantic segmentation (SS‑SS) offers a compelling solution by aiming to achieve strong performance while drastically reducing annotation requirements, by primarily leveraging large quantities of readily available unlabeled data.
Among various SS-SS strategies, Consistency Regularization (CR) has proven highly effective and is widely adopted, often implemented within teacher-student frameworks [8, 9]. In this setup, a “teacher” model (e.g. an exponential moving average (EMA) of the “student”) generates pseudo-labels on weakly perturbed unlabeled images, which then supervise the student model trained on strongly perturbed versions of the same images. This encourages the model to produce consistent predictions despite input perturbations. While CR methods have significantly boosted overall segmentation accuracy (e.g. mIoU), they typically focus on pixel-wise classification without explicit boundary modeling. Notably, recent fully supervised methods have shown that incorporating boundary detection as an auxiliary multi-task learning (MTL) improves segmentation quality, particularly at object boundaries [10, 11, 12]. These boundary-aware approaches help delineate objects more precisely—critical for applications like autonomous driving where clear object separation affects safety.
@@
While some SS-SS methods have begun exploring boundary information, existing approaches typically fall into two categories: (1) methods that derive boundary cues from segmentation pseudo-labels to re-weight consistency losses near edges [13, 14], and (2) approaches like BoundaryMatch [15] that add binary boundary detection as an auxiliary task. However, both categories have limitations. The former relies on boundaries derived from potentially noisy pseudo-labels, which can propagate segmentation errors to boundary regions. The latter, while adding boundary detection as an auxiliary task, still derives boundary pseudo-labels from segmentation outputs via Laplacian operators. This means both tasks are supervised by the same underlying signal, potentially limiting the benefits of multi-task learning where independent supervision signals could provide complementary information. This analysis motivates our approach: learning semantic boundaries from hierarchical features independently of segmentation predictions, while establishing explicit fusion pathways that enable bidirectional information exchange between tasks. This design aims to provide more independent supervision while maintaining beneficial interactions between the two tasks.
Building on this insight, we propose BoundMatch, a teacher-student framework that integrates semantic boundary detection with consistency regularization for SS-SS. At its core, BoundMatch introduces Boundary Consistency Regularized Multi-Task Learning (BCRM), which enforces prediction agreement between teacher and student models on both segmentation masks and semantic boundaries. Crucially, our boundary head learns from hierarchical backbone features through independent task heads, rather than deriving boundaries from segmentation outputs, enabling it to provide complementary supervision. To leverage this complementarity, we incorporate two fusion modules: Boundary-Semantic Fusion (BSF) integrates learned boundary cues into the segmentation decoder, while Spatial Gradient Fusion (SGF) refines boundary predictions using segmentation gradients, yielding cleaner boundary pseudo-labels for consistency regularization. As a foundation for our experiments, we also develop SAMTH, a strong baseline that addresses training instabilities in teacher-student frameworks through Harmonious Batch Normalization (HBN), which maintains consistent batch normalization statistics between teacher and student networks. We demonstrate that BoundMatch’s modular design can enhance various existing CR methods by integrating it with multiple baselines (i.e. SAMTH, UniMatch, and PrevMatch).
Our experiments demonstrate that BoundMatch consistently improves performance across diverse evaluation settings. On urban-driving datasets (Cityscapes, BDD100K, and SYNTHIA) under varying label ratios, BoundMatch achieves improvements of 0.4–2.4% mIoU over strong SAMTH baseline, with state-of-the-art results on the recent DINOv2-based Cityscapes benchmark. The method also performs well in realistic large-scale scenarios with abundant unlabeled data and generalizes to academic benchmarks (Pascal VOC and ADE20K) and remote-sensing data (LoveDA), suggesting that boundary-aware learning can provide benefits across different scenarios. Through ablation studies, we analyze each component’s contribution and demonstrate that improvements extend beyond mIoU to boundary-specific metrics (BIoU and BF1), providing a more complete picture of segmentation quality. Furthermore, BoundMatch scales to lightweight architectures (MobileNet-V2, AFFormer), maintaining its benefits while meeting real-time deployment constraints.
@@
  • We investigate whether semantic boundaries can provide additional supervision in consistency regularization for SS-SS. Through experiments with BCRM, we show that enforcing consistency on both segmentation and boundary predictions from hierarchical features can provide improvements.
  • We design BoundMatch with two fusion modules (BSF and SGF) that connect boundary and segmentation tasks. Our ablations show these modules contribute incrementally to performance, with the complete framework further improving both mIoU and boundary metrics.
  • We identify training instabilities in teacher-student frameworks and propose Harmonious Batch Normalization (HBN) as a solution. While orthogonal to our boundary work, HBN improves several baseline methods and yields SAMTH, which serves as a strong baseline for our experiments.
  • We report boundary-specific metrics (BIoU and BF1) alongside mIoU in our experiments, providing additional perspective on segmentation quality. Our results show that improvements extend to these boundary metrics, though we acknowledge that comprehensive evaluation of all prior methods remains infeasible.
  • We demonstrate that BoundMatch can be integrated with different baseline methods (UniMatch and PrevMatch) and architectures (ResNet, DINOv2-based transformers, lightweight models), showing consistent though incremental improvements across these variations.
@@ 2 Related Work
Table 1: Comparison of boundary-aware semi-supervised segmentation methods. Our BoundMatch utilizes semantic boundaries to further discriminate between different classes, while prior methods rely on binary edges. Unlike previous approaches that derive boundaries from segmentation outputs (i.e. pseudo-labels), BoundMatch learns boundaries independently using a teacher model. Furthermore, we introduce explicit fusion mechanisms (BSF and SGF) to create bidirectional information flow between boundary and segmentation tasks, rather than treating them as independent objectives. These design choices enable BoundMatch to achieve performance gains, evaluated using both standard mIoU and boundary-specific metrics (BIoU and BF1).
CFCG CW-BASS BoundaryMatch BoundMatch (Ours)
2023 [13] 2025 [14] 2024 [15]
Boundary Type Binary Binary Binary Semantic
Boundary Source Derived (Laplacian) Derived (Sobel) Derived (Laplacian) Learned (Teacher pred.)
Integration Loss reweighting Loss reweighting Parallel MTL Fused MTL (BSF+SGF)
Boundary Consistency Reg.
Bidirectional Flow
Feature Stages Single-stage Hierarchical
@@
Fully supervised semantic segmentation methods require large amounts of costly pixel-wise annotations. SS-SS aims to mitigate this burden by leveraging abundant unlabeled data alongside a smaller labeled set. Consistency regularization (CR) has become mainstream in SS-SS, aiming to produce more reliable pseudo-labels and to manage noisy pseudo-labels via perturbation agreement [30, 31, 32, 33, 34, 35, 36]; analyzed by NRCR [37] in terms of label refinement, strong regularization, multi-view learning, robust losses, and sample selection. A common implementation involves the mean-teacher framework [8], where an exponential moving average (EMA) of the student model’s weights generates stable pseudo-labels for unlabeled data. CutMix-MT [8] integrated CutMix augmentation to enhance consistency. Subsequent works like AugSeg [38] and iMAS [39] explored stronger augmentation strategies. Teacher-Student framework has been further extended with various techniques, such as using multiple teachers (e.g. DualTeacher [40]) and contrastive learning (e.g. ReCo [41], U2PL [42]).
There are other notable directions in CR. Some methods focus on co-training to provide cross supervision by using two or more independent networks with different initializations or architectures (e.g. CPS [43] and Diverse Co-Training [44]). On the other hand, approaches like UniMatch [9] and its variants [45, 46] have proposed a unified framework that combines strong-to-weak perturbation regularization with a single model training strategy.
@@
While most SS-SS methods concentrate on pseudo-label refinement and consistency losses, explicit integration of boundary information remains relatively unexplored. Tab.˜1 summarizes the key differences among existing boundary-aware SS-SS approaches.
Some methods derive boundary masks from segmentation pseudo-labels using edge detection operators: CFCG [13] applies Laplacian operators while CW-BASS [14] uses Sobel filters to identify boundary regions for loss re-weighting. These approaches are inherently limited to the boundary information already present in the pseudo-labels.
Other methods explicitly learn boundaries as an auxiliary task. Graph-BAS3Net [5] pioneered this direction for semi-supervised medical image segmentation. BoundaryMatch [15] recently added binary boundary detection to UniMatch through parallel multi-task learning. However, as shown in Tab.˜1, BoundaryMatch derives its boundary pseudo-labels from segmentation outputs using Laplacian operators, potentially propagating noise from the segmentation predictions. Additionally, it uses only binary edges and single-stage features, reporting limited gains when integrated with modern SS-SS methods.
Our BoundMatch explores a different combination of design choices: (1) Semantic boundaries rather than binary edges, providing class-specific information; (2) Learned boundary predictions from a teacher model rather than deriving them from pseudo-labels; (3) Bidirectional fusion between tasks (BSF and SGF) rather than parallel multi-task learning; (4) Hierarchical features from multiple backbone stages [27, 11] for multi-scale boundary reasoning. Table 1 summarizes the key differences among existing boundary-aware SS-SS approaches.
These design choices enable BoundMatch to achieve performance gains as measured by both standard mIoU and boundary-specific metrics (BIoU and BF1). Our results suggest that integrating learned semantic boundaries with explicit fusion mechanisms between boundary and segmentation tasks can be more effective than treating them as independent objectives.
@@ 3 Approach
In this section, we present BoundMatch, our multi-task framework that incorporates boundary detection into semi-supervised semantic segmentation (SS-SS). Our design integrates multiple noise-handling mechanisms identified in recent SS-SS literature [37]: EMA-based label refinement, strong data augmentation, multi-task learning with independent segmentation and boundary heads, and confidence-based sample selection. The boundary detection task, learned from hierarchical features rather than derived from segmentation outputs, provides complementary supervision that can be particularly beneficial at object boundaries.
We begin by outlining the notation and preliminaries for SS-SS, including the teacher-student paradigm, in Secs.˜3.1 and 3.2. We then introduce SAMTH, our baseline method detailed in Sec.˜3.3. Building upon this baseline, Sec.˜3.4 describes our core Boundary Consistency Regularized Multi-Task Learning (BCRM) framework. Subsequently, Sec.˜3.5 details the complementary boundary-aware modeling and refinement strategies: Boundary-Semantic Fusion (BSF) and Spatial Gradient Fusion (SGF). Finally, Sec.˜3.6 synthesizes these components (SAMTH, BCRM, BSF, and SGF) into the complete BoundMatch framework.
@@ 3.2 Teacher-Student Framework for Consistency Regularization
The teacher-student framework, as shown in Fig.˜2, is a popular paradigm for SS-SS. It employs consistency regularization (CR) between the predictions of a teacher network and a student network for unlabeled samples. The core idea is to encourage the student network to produce consistent predictions for an unlabeled image under different perturbations. The teacher network, which is typically an exponential moving average (EMA) of the student network, provides more stable and reliable predictions, which are then used to guide the student’s learning.
@@ 3.3 Strong Data Augmentation Mean-Teacher with Hard Pseudo-Labels (SAMTH)
HBN aims to mitigate these concerns. We forward the complete batch ($x,u^{w},u^{s}$) through both networks: $p^{S,L},p^{S,w},p^{S,s}=f_{\theta^{S}}(x,u^{w},u^{s})$ and $p^{T,L},p^{T,w},p^{T,s}=f_{\theta^{T}}(x,u^{w},u^{s})$. Crucially, the teacher’s BN layers remain in train mode, updating statistics independently based on the full data stream passing through the teacher’s own parameters. For every iteration, the teacher updates its BN statistics, while the rest of the model parameters (i.e. layer weights) are EMA-updated from the student’s parameters. This ensures the teacher’s BN statistics are harmonious with its specific weights and the full input distribution, decoupling its normalization from student fluctuations. This promotes more stable training dynamics and improves performance, as validated empirically across multiple baselines in our ablations (Sec.˜4.3.7). For more complete details and pseudocode, please refer to Appendix A.